# New York, US

### 05 Jan 2024 v1

@@ IV-A RBF Recursive Markov Model
Our Markov Model for the dynamics of the RBF has $\sigma+1$ states labeled $0,1,\cdots,\sigma$. The state labels correspond to the number of bits set to $1$ in the RBF for a given cycle. After the arrival of a unique message (i.e., not yet seen this cycle), depending on the outcome of the hash operation, the number of set bits in the RBF may or may not change. We represent this event by transition probabilities between the states $\tau_{k}{i,j}$. The parameter $k$ determines which transition probabilities are nonzero– e.g. from state $i$, the “largest” possible state we can transition to is state $\text{min}(i+k,\sigma)$. A partial diagram of our Markov Model, displaying only the “forwards” transition behavior, is shown in Figure 1.
@@
We model the resetting of the RBF with “backwards” transition probabilities from certain states $j$ to earlier states $i<j$. In practice, RBFs can implement its recycling behavior in two ways: if message $m$ causes the RBF to reset ($b>\sigma)$, after clearing all the bits, we can:
@@
We call (i) the retaining version of the RBF and (ii) the nonretaining version. For both the retaining and non-retaining RBF, nonzero backwards transition probabilities exist at states $\{\sigma-k+1,\sigma-k+2,...,\sigma\}$. For the retaining RBF, the backwards transitions are to states $\{1,2,...,k\}$. For the non-retaining RBF, all backwards transitions are to state 00. While we analyze both versions, the non-retaining version is slightly simpler for analysis purposes and is the version we work with unless otherwise mentioned. Figure 2 shows a partial diagram of our Markov Model displaying the backward transitions.
Figure 2: Partial Markov Chain diagram for Recycling Bloom Filter showing the “backwards” transition behavior, for $k=2$ (some transition arrows omitted). The light grey arrow depicts the transition for the retaining version, while the dark grey arrow is for the non-retaining version.
@@ IV-C RBF Long-Term Average False Positive Rate
With our Markov Model specified, we can derive an expression for the long-term average False Positive Rate of the RBF, for each combination of colliding/non-colliding and retaining/non-retaining versions (four in total).
@@ IV-C1 Transition Probabilities
We start by deriving expressions for the transition probabilities $\tau_{k}(i,j)$ of an arriving (new message), where $k$ is the number of hash functions, $i$ is the number of bits set to 1 in the BF prior to insertion of the arriving message, and $j$ is the ending number of bits set to 1 (after the next message is hashed into the filter). The definition is agnostic to recycling, such that $i>j\rightarrow\tau_{k}(i,j)=0$. We accomplish this by recursively defining $\tau_{k}(i,j)$ in terms of $\tau_{k-1}(i,j)$ and $\tau_{k-1}(i,j-1)$. Note that the explicit recursive equation depends upon whether the BF is being implemented using colliding (vs. non-colliding) hash functions, as well as whether it is retaining (vs. non-retaining). The intuition behind the recursion is that the $k$th hash function will add at most one bit in the BF after the other $k-1$ hashes have been applied. The BF transitions from $i$ to $j$ bits being set if either a) the first $k-1$ hashes transition to $j$ bits and the last hash is to a bit previously set (by earlier messages and/or the preceding $k-1$ hashes of the current message), or b) the first $k-1$ hashes transition to $j-1$ bits set, and the last hash maps to an unset bit. Note further that these recursive equations make use of the independence properties of the hash functions applied within Bloom Filters, namely that each messages’ hashes are independent of those of previous hashes, that colliding hashes for a given message are also independent from one another, and that non-colliding hashes are chosen via sampling without replacement. The former property ensures that a message arriving to a BF with $b$ bits set can assume these $b$ bits were simply sampled uniformly at random.
@@ IV-C2 Steady-state Probabilities
We can solve the equations for all $i$ by starting with an arbitrary positive assignment (a “guess”) for $\pi_{0}$; call this $g_{0}$. Then, we apply (12) to recursivley compute $g_{i}$ for $\pi_{i},i>0$ as functions of $g_{0}$. Since the sum over all steady state probabilities must be 1, we can renormalize each $\pi_{i}=\frac{g_{i}}{\sum_{j=0}^{\sigma}g_{j}}$.
@@ IV-C3 Expression for Long-Term Average False Positive Rate
Then overall long-term average False Positive rate for the RBF is given by (using the desired definition of $\rho(i)$ from (17):
@@
IV-D Closed-Form expression for one hash function
@@
$f_{\sigma}^{i}=\sum_{i=0}^{\sigma}\frac{i}{M(M-i)\sum_{j=0}^{\sigma}\frac{1}{M-j}}$ (19)
@@ IV-F Expected messages within a $\sigma$-bounded RBF
In each cycle of $\sigma$-bounded RBF, the number of messages that are hashed into the BF prior to recycling will vary, depending on the number of hash collisions between (and for colliding, within) messages. We conclude our analysis of the one-phase BF by showing how the expected number of messages can be computed.
@@
where $X_{j}(i)$ is an indicator that equals 1 only when $j$ additional bits get set from arrival of the $i$th message.
Noting the above equation holds irrespective of the value of $i$, we can simply replace $b(i)$ with $b$ and just write $N(b)$, which we replace with $N_{b}$ to get the result $N_{b}=1+\sum_{j=0}^{k}X_{j}(i)N_{b+j}$. We can then substitute in the r.v. $X_{b,b+j}$ for $X_{j}(i)$ which indicates that this $i+1$st message takes the BF from having $b$ bits set to $b+j$. This can be solved to permit a reverse recursion:
@@ IV-F1 Two-Phase RBF Long-term average False Positive Rate
We can use the results from the standard RBF to also derive an expression for the False Positive rate of the two-phase RBF variant. Let $F_{i}$ denote the steady-state probability of the frozen filter having $i$ bits. Note $F_{i}$ is only nonzero for $i$ between $\sigma-k+1$ and $\sigma$. To compute $F_{i}$, we first define an un-conditional (unnormlized) value of $\hat{F}_{i}$ to then compute a normalized (conditional) version of $F_{i}$ in terms of $\hat{F}_{i}$:
@@ V Results
    We verify the accuracy of our models through discrete event-driven simulations. Simulations were implemented in Python. The RBF data structures were implemented using standard Python libraries. The uniform hash function generation and random message arrival process leveraged the Python random library.
@@
    With the accuracy of the models verified, we turn to the question of the maximum message capacity achievable by a RBF while staying below a given False Positive rate. We consider three models of a RBF: the $\sigma$-bounded model, the lower-bound $f_{a}$, and the traditional equation of (1), $f_{w}$. The $\sigma$-bounded model predicts the highest number of messages, followed by $f_{o}$, $f_{a}$, and $f_{w}$ the lowest. This sequence implies that $N$-bounding variants are overly conservative estimates. Therefore, if one aims to size a RBF to achieve the best performance, the $\sigma$-bounded would serve as the preferred guide for sizing the parameter.
@@
    Finally, for the $\sigma$-bounded model, we investigate the “tradeoff” between using a one-phase and two-phase RBF variant, in terms of the maximum memory capacity achievable by each variant while staying below a given False Positive rate.
@@ V-A1 One and Two-Phase Sigma Bound
We verify the accuracy of $f_{\sigma}^{1}$ and $f_{\sigma}^{2}$ through discrete event-driven simulations. For all simulations, we have a fixed RBF size of $M=1000$ bits. The two-phase RBF splits this memory evenly between each filter. We run multiple simulation epochs of $100,000$ message arrivals, recording the average False Positive rate at the end of each epoch. Message arrivals are drawn uniformly from a pre-generated distribution of $1000$ messages. Given a collection of average False Positive rates, we can plot sample means and confidence intervals using standard statistical methods [24]. We compare these results to the values predicted by $f_{\sigma}^{1}$ and $f_{\sigma}^{2}$. Figure 3 depicts the results for $f_{\sigma}^{1}$. We simulated for a total of $7$ epochs and the $99$% confidence interval is indicated on the plot by the shaded area around the simulation curve. For all data points, the value of $f_{\sigma}^{1}$ lies within the confidence interval. This is strong evidence in favor of the accuracy of $f_{\sigma}^{1}$in modelling False Positive rates for the one-phase RBF.
Similarly, Figure 4 shows results for $f_{\sigma}^{2}$. We simulated for a total of $10$ epochs and the $99$% confidence interval is indicated on the plot. The value of $f_{\sigma}^{2}$ lies within the confidence interval for all points, strong evidence in favor of the accuracy of the model for the two-phase RBF.
Figure 3: One-Phase RBF: verifying our model accuracy through simulation.
Figure 4: Two-Phase RBF: verifying our model accuracy through simulation.
@@ V-A2 Lower bounds on False Positive rate
We turn next to verifying the accuracy of $f_{a}$ and $f_{o}$. For this, we ran $14$ simulation epochs of $1,000,000$ messages each. Figure 5 shows the simulation results plotted against $f_{a}$ and $f_{o}$, with the $99$% confidence interval shown. For all data points, both $f_{o}$ and $f_{a}$ are below the sample mean, and fall either within or below the confidence interval. This demonstrates strong evidence in favor of the accuracy of the lower bounds $f_{o},f_{a}$. Observe that $f_{a}$ is a “tighter” lower lower bound than $f_{o}$.
@@ V-B Expected Message Capacity Comparison
A natural question to ask when optimizing RBF parameters is the expected maximum number of messages one can store while staying under an acceptable average False Positive rate– we refer to this as the expected message capacity. For the one-phase $\sigma$-bounded model (False Positive rate given by $f_{\sigma}^{1}$), recall this is $E[N_{0}]$ derived in §IV-F. We can also compute an expected message capacity using $f_{a}$ or $f_{\omega}$. Figure 6 compares the expected message capacity between the three models. For each model, we fix an average False Positive rate of .01.01.01.01. For each value of $M$, we find the maximum number of messages subject to the False Positive constraint. We plot this number normalized to $f_{\sigma}^{1}$, as this is the model that yields the highest message capacity. $f_{a}$ is seen to be a good approximation, a direct consequence of the tightness of the lower bounds discussed in §III. Observe that $f_{\omega}$ can lead to an overly conservative estimate of message capacity in this context; for the given parameters it consistently suffers a $>30$% underestimation of maximum message capacity.
Figure 7. shows similar results, this time varying False Positive rates on the x-axis. Once again, the worst-case $N$ model sees a consistent underestimation of max message capacity of around $30$%.
@@ V-C Tradeoff between One and Two-phased RBF
We can also consider the question of which is the “better performing” of $f_{\sigma}^{1}$ or $f_{\sigma}^{2}$ when it comes to average False Positive rate.
In Figure 8, we make this comparison. We have a given filter memory $M$ (which in the case of the two-phase RBF is split evenly between both filters) and False Positive rate. We then find the $k$ value which allows us the maximum number of expected messages for these constraints. In both cases, we observe the result that the one-phase RBF appears to outperform the two-phased filter when it comes to squeezing out extra filter bit capacity. One might be moved to ponder what is the point of the added complexity of a two-phased filter, if it apparently is “less efficient” in terms of average False Positive rate? The answer becomes apparent when we consider the idea of false negatives as we previously mentioned. While a detailed investigation of false negatives is beyond the scope of this paper, the intuition is that the extra bit capacity seemingly afforded by the one-phased filter comes with an increased probability of false negative rates, which are mitigated by the two-phased filter.
@@ VI Related Work
Luo et al. provides a general summary of Bloom Filters and their variants, along with their applications to problems in computing [14]. To the best of our knowledge, all previous analyses of Bloom Filters approach the question of False Positive rates from the perspective of a “worst case” bound; i.e., the False Positive rate of an $n+1$st message after $n$ messages have been received. Equation (1) is widely used in most applications to compute the False Positive rate according to this metric. Two prior works have shown this equation is slightly inaccurate, and in fact a lower bound for the true “worst case” False Positive rate [4, 5].
RBFs have been employed in the classification of known/authorized traffic within the data plane of Software-Defined Networks (SDN) [19], in the detection of DDoS attacks [6], stateful load balancing to ensure per-connection consistency in datacenters [20] and payload attribution [23]. In all these scenarios, the traditional “worst case” False Positive analysis may be less useful. For instance, the worst case analysis may yield a much stricter than necessary bound on the maximum number of BF message insertions. The False Positive analysis presented in this paper offers an alternative, and in some cases more practical approach to designing BF parameters for these applications.
There have been other BF variants proposed that consider the need to periodically remove items from the filter. The Counting Bloom Filter (CBF) [10] and similar variants support deletion of individual elements by replacing the bits of the BF with $l$-bit counters. These counters are incremented or decremented upon message insertion or deletion. CBF’s are utilized as a way to support element removal in applications such as [28], [21]. However, the CBF and its variants all come with increased space and algorithmic overhead to support deletion in the first place. Among similar lines, variants such as the Deletable Bloom Filter [22], the Ternary Bloom Filter [13] and the Quotient Filter [2] support element deletion under restricted circumstances, also at the cost of increased space and algorithmic complexity. The idea of periodically resetting a targeted subset of BF bits is first proposed in Donnet et al. [7]; while this method is shown to reduce the overall False Positive rate, it also comes at the cost of increased algorithmic complexity in determining the selected subset of bits. More recently, Cuckoo Filters have been used as an extension of Bloom filters. Cuckoo filters use Cuckoo hashing, a technique applied for resolving hash table collisions by continually relocating hashed entries until they land on unoccupied bins, to optimize space utilization. These filters notably offer the ability to delete elements post-insertion without practical overhead [9]. A caveat to this, however, is the prerequisite knowledge of the elements that necessitate deletion, a condition not commonly met in many networking applications. In addition, unlike RBFs, Cuckoo filters possess a finite capacity. This implies that once a certain number of elements are added, the filter reaches its maximum capacity and cannot accommodate additional entries until existing ones are removed. This inherently limits the scope of its usage under conditions where an unknown and continuous influx of elements is anticipated, which are common in many networking applications.

### 03 Feb 2024 v2

@@ IV-A RBF Recursive Markov Model
Our Markov Model for the dynamics of the RBF has $\sigma+1$ states labeled $0,1,\cdots,\sigma$. The state labels correspond to the number of bits set to $1$ in the RBF during a given cycle. After the arrival of a unique message (i.e., not yet seen this cycle), depending on the outcome of the hash operation, the number of set bits in the RBF may or may not change. We represent this event by transition probabilities between the states, labeled $\tau_{k}{i,j}$. The number of employed hash functions, $k$, determines which transition probabilities are nonzero– e.g. from state $i$, the “largest” possible state we can transition to is state $\text{min}(i+k,\sigma)$. A partial diagram of our Markov Model, displaying only the “forward” transition behavior, is shown in Fig.  1.
@@
We model the resetting of the RBF with “backwards” transition probabilities from certain states $j$ to earlier states $i<j$. In practice, RBFs can implement recycling behavior in two ways: if message $m$ causes the RBF to reset ($b>\sigma)$, after clearing all the bits, we can:
@@
We call (i) the retaining version of the RBF and (ii) the nonretaining version. For both the retaining and non-retaining RBF, nonzero backwards transition probabilities exist at states $\{\sigma-k+1,\sigma-k+2,...,\sigma\}$. For the retaining RBF, the backwards transitions are to states $\{1,2,...,k\}$. For the non-retaining RBF, all backwards transitions are to state 00. While we analyze both versions, the non-retaining version is slightly simpler for analysis purposes and is the version we work with unless otherwise mentioned. Fig.  2 shows a partial diagram of our Markov Model displaying the backward transitions.
Figure 2: Partial Markov Chain diagram for Recycling Bloom Filter showing the “backwards” transition behavior, for $k=2$ (some transition arrows omitted). The light grey arrow depicts the transition for the retaining version, while the dark grey arrow is for the non-retaining version, non-colliding.
@@ IV-C RBF Long-Term Average False Positive Rate
With our Markov Model specified, we can derive an expression for the long-term average False Positive Rate of the RBF, for both colliding and non-colliding BFs, and also derive the backward transitions for colliding/noncolliding and retaining/non-retaining versions (four in total).
@@ IV-C1 Transition Probabilities
We start by deriving expressions for the forward transition probabilities $\tau_{k}(i,j)$ of an arriving (new message), where $k$ is the number of hash functions, $i$ is the number of bits set to 1 in the BF prior to insertion of the arriving message, and $j$ is the ending number of bits set to 1 (after the next message is hashed into the filter). We accomplish this for forward transitions from $i$ to $j\geq i$ by recursively defining $\tau_{k}(i,j)$ in terms of $\tau_{k-1}(i,j)$ and $\tau_{k-1}(i,j-1)$. The explicit recursive equation depends upon whether the BF is being implemented using colliding (vs. non-colliding) hash functions, as well as whether it is retaining (vs. non-retaining). The intuition behind the recursion is that the $k$th hash function will add at most one bit in the BF after the other $k-1$ hashes have been applied. The BF transitions from $i$ to $j$ bits being set if either a) the first $k-1$ hashes transition to $j$ bits and the last hash is to a bit previously set (by earlier messages and/or the preceding $k-1$ hashes of the current message), or b) the first $k-1$ hashes transition to $j-1$ bits set, and the last hash maps to an unset bit. Note that these recursive equations make use of the independence properties of the hash functions applied within Bloom Filters; each message’s hashes are independent of those of previous hashes, colliding hashes for a given message are also independent from one another, and non-colliding hashes are chosen via sampling without replacement. The former property ensures that a message arriving to a BF with $b$ bits set can assume these $b$ bits were simply sampled uniformly at random.
@@ IV-C2 Steady-state Probabilities
We can solve the equations for all $i$ by starting with an arbitrary positive assignment (a “guess”) for $\pi_{0}$; call this $g_{0}$. Then, we apply (12) to recursively compute $g_{i}$ for $\pi_{i},i>0$ as functions of $g_{0}$. Since the sum over all steady state probabilities must be 1, we can renormalize each $\pi_{i}=\frac{g_{i}}{\sum_{j=0}^{\sigma}g_{j}}$.
@@ IV-C3 Expression for Long-Term Average False Positive Rate
Then the overall long-term average False Positive rate for the RBF is given by (using the desired definition of $\rho(i)$ from (17):
@@
IV-D Closed-Form expression for $k=1$
@@
$f_{\sigma}^{1,k=1}=\sum_{i=0}^{\sigma}\frac{i}{M(M-i)\sum_{j=0}^{\sigma}\frac{1}{M-j}}$ (19)
@@ IV-F Expected messages within a $\sigma$-bounded RBF
In each cycle of a $\sigma$-bounded RBF, the number of messages that are hashed into the BF prior to recycling will vary, depending on the number of hash collisions between (and for colliding, also within) messages. We conclude our analysis of the one-phase BF by showing how the expected number of messages can be computed.
@@
where $X_{j}(i)$ is an indicator that equals 1 only when $j$ additional bits get set from the arrival of the $i$th message.
Noting the above equation holds irrespective of the value of $i$, we can simply replace $b(i)$ with $b$ and just write $N(b)$, which we replace with $N_{b}$ to get the result $N_{b}=1+\sum_{j=0}^{k}X_{j}(i)N_{b+j}$. We can similarly substitute in the r.v. $X_{b,b+j}$ for $X_{j}(i)$ which indicates that this $i+1$st message takes the BF from having $b$ bits set to $b+j$. This can be solved to permit a reverse recursion:
@@ IV-F1 Two-Phase RBF Long-term average False Positive Rate
We can use the results from the standard RBF to also derive an expression for the False Positive rate of the two-phase RBF variant. Let $F_{i}$ denote the steady-state probability of the frozen filter having $i$ bits. Note $F_{i}$ is only nonzero for $i$ between $\sigma-k+1$ and $\sigma$. To compute $F_{i}$, we first define an un-conditional (non-normalized) value of $\hat{F}_{i}$ to then compute a normalized (conditional) version of $F_{i}$ in terms of $\hat{F}_{i}$:
@@ V Results
    We verify the accuracy of our models through discrete event-driven simulations implemented in Python. The RBF data structures were implemented using standard Python libraries. The uniform hash function generation and random message arrival process leveraged the Python random library. Code can be found at: [1]
@@
    With the accuracy of the models verified, we turn to the question of the maximum message capacity achievable by a RBF while staying below a given False Positive rate. We consider four models of a one-phase RBF: the $\sigma$-bounded model (18), the “oracle” $N$-bounded False Positive rate (2), the lower-bound on the $N$-bounded average False Positive rate, $f_{a}$ (3), and the traditional worst-case bound $f_{w}$ (1). The $\sigma$-bounded model maintains the highest number of messages, followed by $f_{o}$, $f_{a}$, and $f_{w}$ the lowest. This sequence implies that $N$-bounding variants are overly conservative estimates. Therefore, if one aims to size a RBF to achieve the best performance, the $\sigma$-bounded variant should be utilized.
@@
    Finally, for the $\sigma$-bounded model, we investigate the trade-off between using a one-phase and two-phase RBF variant, in terms of the maximum memory capacity achievable by each variant while staying below a given False Positive rate.
@@ V-A1 One and Two-Phase Sigma Bound
We verify the accuracy of $f_{\sigma}^{1}$ and $f_{\sigma}^{2}$ through discrete event-driven simulations. For all simulations, we have a fixed RBF size of $M=1000$ bits. The two-phase RBF splits this memory evenly between each filter. We run multiple simulation epochs of $100,000$ message arrivals, recording the average False Positive rate at the end of each epoch. Message arrivals are drawn uniformly from a pre-generated distribution of $1000$ messages. Given a collection of average False Positive rates, we can plot sample means and confidence intervals using standard statistical methods [18]. We compare these results to the values predicted by $f_{\sigma}^{1}$ and $f_{\sigma}^{2}$. Fig. 3 depicts the results. We simulated for a total of $7$ epochs for $f_{\sigma}^{1}$ and $10$ epochs for $f_{\sigma}^{2}$. The $99$% confidence interval is indicated on the plot by the shaded area around the simulation curves. For all data points, the value of $f_{\sigma}^{1}$ and $f_{\sigma}^{2}$ lies within the confidence interval. This is strong evidence in favor of the accuracy of $f_{\sigma}^{1}$ and $f_{\sigma}^{2}$ in modelling False Positive rates for RBFs.
Figure 3: Verifying our model accuracy through simulation.
@@ V-A2 Lower bounds on False Positive rate
We turn next to verifying the accuracy of $f_{a}$ and $f_{o}$. For this, we ran $14$ simulation epochs of $1,000,000$ messages each. Fig. 4 shows the simulation results plotted against $f_{a}$ and $f_{o}$, with the $99$% confidence interval shown. For all data points, both $f_{o}$ and $f_{a}$ are below the sample mean, and fall either within or below the confidence interval. This demonstrates strong evidence in favor of the accuracy of the lower bounds $f_{o},f_{a}$. Observe that $f_{a}$ is a “tighter” lower lower bound than $f_{o}$.
@@ V-B Expected Message Capacity Comparison
A natural question to ask when optimizing RBF parameters is the expected maximum number of messages one can store while staying under an acceptable average False Positive rate– we refer to this as the expected message capacity. For the one-phase $\sigma$-bounded model (False Positive rate given by $f_{\sigma}^{1}$), recall this is $E[N_{0}]$ derived in §IV-F. We can also compute an expected message capacity using $f_{a}$ or $f_{\omega}$. Fig. 5 compares the expected message capacity between the three models. For each model, we fix an average False Positive rate of .01.01.01.01. For each value of $M$, we find the maximum number of messages subject to the False Positive constraint. We plot this number normalized to $f_{\sigma}^{1}$, as this is the model that yields the highest message capacity. $f_{a}$ is seen to be a good approximation, a direct consequence of the tightness of the lower bounds discussed in §III. Observe that $f_{\omega}$ can lead to an overly conservative estimate of message capacity in this context; for the given parameters its maximum message capacity is consistently reduced by more than 30%.
@@
Fig. 6. shows similar results, this time varying False Positive rates on the x-axis. Once again, the worst-case $N$ model sees a consistent underestimation of maximum message capacity of around $30$%.
@@ V-C Tradeoff between One and Two-phased RBF
In Fig. 7, we compare the False Positive rates of one-phase RBFs ($f_{\sigma}^{1}$) versus their two-phase counterparts ($f_{\sigma}^{2}$). We have a given filter memory $M$ (which in the case of the two-phase RBF is split evenly between both filters) and False Positive rate. We then find the value of $k$ which allows us the maximum number of expected messages for these constraints. In both cases, we observe the result that the one-phase RBF appears to outperform the two-phased filter when it comes to squeezing out extra filter bit capacity. One might be moved to ponder what is the point of the added complexity of a two-phased filter, if it apparently is “less efficient” in terms of average False Positive rate? The answer becomes apparent when we consider the idea of False Negatives as we previously mentioned. While a detailed investigation of False Negatives is beyond the scope of this paper, the intuition is that the extra bit capacity seemingly afforded by the one-phase filter comes with an increased probability of False Negative rates, which are mitigated by the two-phase filter.
@@ VI Related Work
Luo et al. provides a general summary of Bloom Filters and their variants, along with their applications to problems in computing [11]. To the best of our knowledge, all previous analyses of Bloom Filters approach the question of False Positive rates from the perspective of a “worst case” bound. Equation (1) is widely used in most applications to compute the False Positive rate according to this metric. Two prior works have shown this equation is slightly inaccurate, and in fact a lower bound for the true “worst case” False Positive rate [4, 5].
There have been other BF variants proposed that consider the need to periodically remove items from the filter. The Counting Bloom Filter (CBF) [8] and similar variants support deletion of individual elements. However, the CBF and its variants all come with increased space and algorithmic overhead to support deletion. Among similar lines, variants such as the Deletable Bloom Filter [17], the Ternary Bloom Filter [10] and the Quotient Filter [2] support element deletion under restricted circumstances, also at the cost of increased space and algorithmic complexity.
The idea of periodically resetting a targeted subset of BF bits is first proposed in Donnet et al. [6]; while this method is shown to reduce the overall False Positive rate, it also comes at the cost of increased algorithmic complexity. More recently, Cuckoo Filters have been used as an extension of Bloom filters. Cuckoo filters use Cuckoo hashing to optimize space utilization. These filters notably offer the ability to delete elements post-insertion without practical overhead [7]. However, this requires knowledge of the elements that necessitate deletion, a condition not commonly met in many networking applications. In addition, unlike RBFs, Cuckoo filters possess a finite unique message capacity. This inherently limits the scope of its usage under conditions where an unknown and continuous influx of elements is anticipated, which are common in many networking applications.